To produce x units of a religious medal costs C(x)=13x+119. The revenue is R(x)=30x. Both cost and revenue are in dollars.
a. Find the​ break-even quantity.
b. Find the profit from 560 units.
c. Find the number of units that must be produced for a profit of ​$170.

Found 2 solutions by macston, stanbon:
Answer by macston(5194)

.
a. Find the break-even quantity.
Break-even is when cost=revenue:
.
C(x)=R(x)
13x+119=30x
119=17x
7=x
ANSWER a): . The break-even quantity is 7.
.
b. Find the profit from 560 units.
P(x)=profit=revenue - cost; x=560
.
P(x)=R(x)-C(x)
P(x)=30x-(13x+119)
P(x)=30x-13x-119
P(x)=17x-119
P(560)=17(560)-119
P(560)=9520-119
P(560)=9401
ANSWER b): . The profit on 560 units is $9401.
.
c. Find the number of units that must be produced for a profit of $170.
.
P(x)=R(x)-C(x)
P(x)=30x-(13x+119)
P(x)=17x-119
$170=17x-119
289=17x
17=x
ANSWER c): . Seventeen units must be produced for a profit of $170.
.
















Answer by stanbon(75887)

In 1998​, there were 41,175 shopping centers in a certain country. In 2008​, there were 48 comma 48,723.
​(a) Write an equation expressing the number y of shopping centers in terms of the number x of years after 1998.
slope = (48723-41,175)/10 = 7548/10 = 754.8
f(1998) = 41,175
y = 754.8x + 41,175
-----------------------------------
​(b) When will the number of shopping centers reach 60,000​?
Solve::
60,000 = 754.8x + 41,175
----
754.8x = 18825
----
x = 24.94
-----
Ans: year = 1998 + 24.94 = 2023 when rounded up
--------------
Cheers,
Stan H.
-----------